jim and bob are panning for gold in leadville,colorado.the ratio of gold found by jim to gold found by bob is 4:7.if the two men together find a total of 55 pieces of gold,find how many pieces of gold each man found.
(Use the 5-step method and check your work)

Answer by ptaylor(2198)

OK. Now we know that for every 7+4=11 pieces of gold found, Jim finds 4 and Bob finds 7. In other words, Jim finds 4/11 of the gold and Bob finds 7/11 of the gold.
(4/11)*55 =20----amount found by Jim
(7/11)*55=35---amount found by Bob
CK
20/35=4/7
Hope this helps---ptaylor